I need help calculating sum of squares of x given the following data as well as the SSxy

13. Calculate SS.x for the following data. Do this however you need to. 20 10 4 12 13 13 12 14. Calculate SSy for the data in question 13.
Answer #1

13) In this part, we have to find SSxx for the data.

Here n = 6

Formula is

SSxx=\sum x^{2}-\frac{(\sum x)^{2}}{n}

x y x^2
10 4 100
11 5 121
12 6 144
13 4 169
13 7 169
12 2 144
Total 71 28 847

SSxx=\sum x^{2}-\frac{(\sum x)^{2}}{n}=847-\frac{71^{2}}{6} =847-840.1667=6.8333

14) Now in this part, we have to find SSxy

Formula is

SSxy=\sum xy-\frac{\sum x\sum y}{n}

x y xy
10 4 40
11 5 55
12 6 72
13 4 52
13 7 91
12 2 24
Total 71 28 334

SSxy=\sum xy-\frac{\sum x\sum y}{n}=334-\frac{71*28}{6}=334-331.333=2.6667